# HiLight: A Hierarchical Reinforcement Learning Framework with Global Adversarial Guidance for Large-Scale Traffic Signal Control

### v1

@@ 1 Introduction
In recent years, RL has emerged as a promising approach for adaptive TSC[4, 5, 6], owing to its ability to learn optimal policies through interaction with the environment. As illustrated in Figure 1(b), existing RL-based TSC methods are broadly categorized into centralized and multi-agent approaches. Centralized RL treats the entire network as a single decision-maker and effectively optimizes global objectives[7, 8, 9], but suffers from poor scalability due to the exponential growth of joint state and action spaces. Multi-agent RL improves scalability by assigning an agent to each intersection[10, 11, 12, 13], yet often lacks global coordination, leading to misaligned local decisions and reduced system-wide efficiency. Challenge 1: How to develop an RL-based TSC method that simultaneously ensures scalability and global coordination remains unresolved.
(a) Real-world traffic flow variations
(b) Multi-agent vs. Centralized RL
Figure 1: Realistic traffic flow scenarios and RL architectures for TSC
Despite the development of various RL algorithms, the evaluation of TSC methods remains constrained by limitations of existing benchmark scenarios. Widely used scenarios (such as Grid5×5[14], Arterial4×4[15], and Cologne8[16]) are based on either synthetic road networks or simplified real-world maps, and the traffic flows are simulated using default configurations or real data sampled from a narrow time window. Challenge 2: These scenarios fail to capture the complexity and variability of real-world traffic, which can be influenced by diverse conditions such as adverse weather, holiday travel surges, or transitions between peak and off-peak periods, as shown in Figure 1(a).
@@ 3.1 Preliminaries
Intersection refers to a location in the road network where two or more roads cross or merge, typically equipped with traffic signals to regulate vehicle movement. Figure 2 shows a typical example: a standard four-leg intersection (also called a four-arm intersection), where four roads meet, forming a cross shape. Each approach (arm) contains three incoming lanes $l_{in}$ dedicated to left-turn, through, and right-turn movements, and typically three corresponding outgoing lanes $l_{out}$.
@@
A phase is defined as the set of green light indications for all incoming lanes at an intersection. As illustrated in Figure 2, a standard intersection typically adopts eight phases, each governing vehicle movements such as left turns, through movements, and right turns. In some scenarios, a single incoming lane may accommodate multiple movement directions. For example, in the Arterial 4×4 scenario, certain lanes serve both through and right-turning vehicles. Additionally, in accordance with real-world practices, right-turning vehicles are not regulated by traffic signals.
@@
Pressure is a metric used to quantify the imbalance of traffic flow at an intersection, specifically between incoming and outgoing lanes. It is typically defined as the difference between the queue lengths (or vehicle densities) of incoming and outgoing lanes for a specific movement. The total pressure $P$ for a phase is usually calculated as the sum of pressures over all movements controlled by that phase: $P=\sum_{(i,j)}(q_{i}-q_{j})$, where $q_{i}$ is the number of vehicles on the incoming lane $l_{in}^{i}$, and $q_{j}$ is that on the corresponding outgoing lane $l_{out}^{j}$, as shown in Figure 2.
@@ 3.2 Problem Formulation
We model the large-scale TSC problem based on Multi-agent Markov Decision Process (MMDPs) as hierarchical MMDPs, which can be defined by the tuple $<\mathcal{I,S,O,A,P,R,\gamma}>$. In this paper, we refer to the high-level policy of HRL as the Meta-Policy, which is responsible for generating abstract goals or selecting sub-policies. The low-level policy, referred to as the Sub-Policy, executes concrete actions to fulfill the objectives specified by the Meta-Policy. The set of intersections is defined as =$\mathcal{I}=\{1,2,\dots,N\}$, where $N$ is the total number of intersections and each $i\in\mathcal{I}$ is controlled by an agent. All agents share the same set of parameters. $S$ denotes the global state, and $O$ is a partial observation derived from $S$ via an observation function $f_{obs}:s\rightarrow o$, where $s\in S,o\in O$. The action space is defined as $A=\{A^{H},A^{L}\}$, where $A^{H}$ contains only a single global sub-goal $a^{H}$ assigned by the Meta-Policy, and each Sub-Policy agent selects its local action $A^{L}=\{a^{L}_{i}\}_{i\in\mathcal{I}}$ based on its local observation and the global sub-goal $a^{H}$. $P$ and $\gamma$ denote the transition probability function and the discount factor separatel. The reward function $\mathcal{R}$ is shared among all intersections and consists of a global high-level reward and local low-level rewards, formulated as:
@@ 4 Methodology
In this section, we propose a novel HRL framework HiLight for traffic signal control. As shown in Figure 3, the architecture consists of two branches: a global Meta-Policy and a local Sub-Policy, which work in coordination while being jointly optimized in an adversarial manner to achieve efficient control across large-scale traffic networks.
Figure 3: An overview of HiLight.
@@ 4.1.1 Transformer Encoding Across Subregions
$[\mathbf{E}_{G}^{t},\mathbf{E}_{1}^{t},\mathbf{E}_{2}^{t},\ldots,\mathbf{E}_{M}^{t}]={TransformerEncoder}(\mathbf{X}^{t}_{{pos}})$
@@ 4.1.2 LSTM-based Sub-goal Generation
[…]
@@ 4.2.2 Dynamic Neighbor Aggregation
[…]
@@ 4.3.1 Reward and Trajectory
The reward structure optimizes global traffic efficiency and local intersection throughput simultaneously. At each time step $t$, each intersection (agent) $i$ receives a local reward $r_{i}^{t}$ that evaluates its own traffic condition. The reward is calculated as a weighted combination of several key metrics: queue length $ql_{i}^{t}$, average vehicle waiting time $wt_{i}^{t}$, delay time $dt_{i}^{t}$, pressure $ps_{i}^{t}$ and speed score $ss_{i}^{t}$, with detailed definitions in Appendix A. The local reward is then given as:$r_{i}^{t}=-\left(ql_{i}^{t}+wt_{i}^{t}+dt_{i}^{t}+ps_{i}^{t}-ss_{i}^{t}\right)$.
@@ Appendix A Detailed Descriptions and Computation of Observations
stop_car_num Number of stationary vehicles on a lane (e.g., stopped due to red lights or congestion). The number of vehicles with speed $<0.1m/s$, normalized by lane capacity; N-dimensional.
@@ Appendix B Ablation Study
To investigate the contribution of each core component in the proposed HiLight framework, an ablation study is conducted by progressively removing key modules and evaluating the resulting performance during training. As shown in the Episode–Reward curves in Figure 5(a), the complete HiLight model achieves the best overall performance, exhibiting both the fastest convergence speed and the most stable behavior after convergence.
(a) Ablation Study
(b) Pretraining Study
Figure 5: The episode–reward curves for ablation and pretraining analysis of the HiLight
@@
We also examine the impact of pretraining the Meta-Policy prior to full model training, as shown in Figure 5(b). The Episode–Reward curve shows a smoother and faster start, without the initial reward drop observed in the baseline. Pretraining allows the Meta-Policy to generate more informative sub-goals in early training, which effectively guides the Sub-Policy toward better optimization directions and improves exploration efficiency. Although the learning slope becomes gentler in later stages, convergence speed is slightly improved, and the final performance remains on par with the non-pretrained HiLight.
@@ Appendix D Visualizations of the scenarios
Figure 6, Figure 7, and Figure 8 illustrate the traffic scenarios used in our experiments.
Figure 6: Synthetic scenarios with sample intersections.
Figure 7: German real-world scenarios with sample intersections
Figure 8: Manhattan real-world scenario with sample intersections
@@ Appendix E Hyper-parameter Settings
Entropy Coefficient $0.01$ Number of Actions $8$
@@
Discount Factor $0.99$ Yellow Light Duration (s) $5$
@@
Learning Rate $43e-$ 4 SUMO Start Port $15000$
@@ Appendix H Broad Impacts
Meta-Policy Transformer Encoder Extract global and regional spatiotemporal features across subregions Transformer Encoder: 3 layers, 2 heads, hidden size 4, feedforward dim = 165, no layer norm $(20,4,4)$: 20 time steps, 4 regions, 4-dim observation per region Adds a class token and positional encoding $\rightarrow$ Transformer encoder $\rightarrow$ output $(5,20,4)$rightarrow extract $(1,1,4)$ as global feature; transform remaining $(4,20,4)$rightarrow $(1,20,16)$
Meta-Policy MultiLayer LSTM Capture global temporal patterns across all regions 4-layer LSTM + FC layer $(1,20,16)$: sequence of flattened regional features LSTM (hidden size = 256) $\rightarrow$ use last-step output $\rightarrow$ FC layer $\rightarrow$ $(1,1,16)$rightarrow reshape to $(1,4,4)$
@@
Feature Fusion Combine global and local features as input for shared policy network Broadcast + Concatenation Global: $(1,1,4)$, Local: $(1,16,330)$ Broadcast global feature to $(1,16,4)$rightarrow concatenate with local features $\rightarrow$ fused tensor $(1,16,334)$
Shared Head for Critic & Actor Project fused traffic features into shared latent representation FC layers: $334\rightarrow 256\rightarrow 128\rightarrow 114$ (with dropout) $(1,16,334)$: concatenated per-node features Three-layer MLP with ReLU and dropout $\rightarrow$ output $(1,16,114)$rightarrow used by actor and critic
@@
Actor Model Output traffic signal control actions 2-layer MLP with ReLU activation $(16,114)$: shared feature per signal $114\rightarrow 57\rightarrow 8\Rightarrow$ action logits (16,)(16,)( 16 , )

### v2

@@ 1 Introduction
In recent years, RL has emerged as a promising approach for adaptive TSC[4, 5, 6], owing to its ability to learn optimal policies through interaction with the environment. As illustrated in Figure, existing RL-based TSC methods are broadly categorized into centralized and multi-agent approaches. Centralized RL treats the entire network as a single decision-maker and effectively optimizes global objectives[7, 8, 9], but suffers from poor scalability due to the exponential growth of joint state and action spaces. Multi-agent RL improves scalability by assigning an agent to each intersection[10, 11, 12, 13], yet often lacks global coordination, leading to misaligned local decisions and reduced system-wide efficiency. Challenge 1: How to develop an RL-based TSC method that simultaneously ensures scalability and global coordination remains unresolved.
Despite the development of various RL algorithms, the evaluation of TSC methods remains constrained by limitations of existing benchmark scenarios. Widely used scenarios (such as Grid5×5[14], Arterial4×4[15], and Cologne8[16]) are based on either synthetic road networks or simplified real-world maps, and the traffic flows are simulated using default configurations or real data sampled from a narrow time window. Challenge 2: These scenarios fail to capture the complexity and variability of real-world traffic, which can be influenced by diverse conditions such as adverse weather, holiday travel surges, or transitions between peak and off-peak periods, as shown in Figure.
@@ 3.1 Preliminaries
Intersection refers to a location in the road network where two or more roads cross or merge, typically equipped with traffic signals to regulate vehicle movement. Figure shows a typical example: a standard four-leg intersection (also called a four-arm intersection), where four roads meet, forming a cross shape. Each approach (arm) contains three incoming lanes $l_{in}$ dedicated to left-turn, through, and right-turn movements, and typically three corresponding outgoing lanes $l_{out}$.
@@
A phase is defined as the set of green light indications for all incoming lanes at an intersection. As illustrated in Figure, a standard intersection typically adopts eight phases, each governing vehicle movements such as left turns, through movements, and right turns. In some scenarios, a single incoming lane may accommodate multiple movement directions. For example, in the Arterial 4×4 scenario, certain lanes serve both through and right-turning vehicles. Additionally, in accordance with real-world practices, right-turning vehicles are not regulated by traffic signals.
@@
Pressure is a metric used to quantify the imbalance of traffic flow at an intersection, specifically between incoming and outgoing lanes. It is typically defined as the difference between the queue lengths (or vehicle densities) of incoming and outgoing lanes for a specific movement. The total pressure $P$ for a phase is usually calculated as the sum of pressures over all movements controlled by that phase: $P=\sum_{(i,j)}(q_{i}-q_{j})$, where $q_{i}$ is the number of vehicles on the incoming lane $l_{in}^{i}$, and $q_{j}$ is that on the corresponding outgoing lane $l_{out}^{j}$, as shown in Figure.
@@ 3.2 Problem Formulation
We model the large-scale TSC problem based on Multi-agent Markov Decision Process (MMDPs) as hierarchical MMDPs, which can be defined by the tuple $\mathcal{I,S,O,A,P,R,\gamma}>$. In this paper, we refer to the high-level policy of HRL as the Meta-Policy, which is responsible for generating abstract goals or selecting sub-policies. The low-level policy, referred to as the Sub-Policy, executes concrete actions to fulfill the objectives specified by the Meta-Policy. The set of intersections is defined as $\mathcal{I}=\{1,2,\dots,N\}$, where $N$ is the total number of intersections and each $i\in\mathcal{I}$ is controlled by an agent. All agents share the same set of parameters. $S$ denotes the global state, and $O$ is a partial observation derived from $S$ via an observation function $f_{obs}:s\rightarrow o$, where $s\in S,o\in O$. The action space is defined as $A=\{A^{H},A^{L}\}$, where $A^{H}$ contains only a single global sub-goal $a^{H}$ assigned by the Meta-Policy, and each Sub-Policy agent selects its local action $A^{L}=\{a^{L}_{i}\}_{i\in\mathcal{I}}$ based on its local observation and the global sub-goal $a^{H}$. $P$ and $\gamma$ denote the transition probability function and the discount factor separatel. The reward function $\mathcal{R}$ is shared among all intersections and consists of a global high-level reward and local low-level rewards, formulated as:
@@ 4 Methodology
In this section, we propose a novel HRL framework HiLight for traffic signal control. As shown in Figure, the architecture consists of two branches: a global Meta-Policy and a local Sub-Policy, which work in coordination while being jointly optimized in an adversarial manner to achieve efficient control across large-scale traffic networks.
@@ 4.1.1 Transformer Encoding Across Subregions
$\mathbf{E}_{G}^{t},\mathbf{E}_{1}^{t},\mathbf{E}_{2}^{t},\ldots,\mathbf{E}_{M}^{t}]={TransformerEncoder}(\mathbf{X}^{t}_{{pos}})$
@@ 4.1.2 LSTM-based Sub-goal Generation
where $\mathbf{h}_{z_{i}}^{T}$ denotes the final hidden state of the LSTM for subregion $z_{i}$, $\|\,$ represents concatenation, and $\phi(\cdot)$ is a feedforward layer to project the concatenated features into a fixed-size global sub-goal vector $\mathbf{G}$. This vector $\mathbf{G}$ is broadcasted to all agents of Sub-Policy and serves as a shared directional objective, encouraging the agents at each intersection to align their local decisions with global traffic optimization trends.
@@ 4.2.2 Dynamic Neighbor Aggregation
with $\mathbf{a}\in\mathbb{R}^{2F}$ a learnable attention vector and $\|\,$ denoting vector concatenation, and attention scores are computed as $\alpha_{ij}=softmax(e_{ij})$ over the neighborhood $\mathcal{N}_{i}=\{j\mid A_{ij}=1\}$. The final representation $\mathbf{z}_{i}^{t}$ is constructed by concatenating $\mathbf{h}_{i}^{t}$ with the weighted neighbor features:
@@ 4.3.1 Reward and Trajectory
[…]
@@ Appendix A Detailed Descriptions and Computation of Observations
stop_car_num Number of stationary vehicles on a lane (e.g., stopped due to red lights or congestion). The number of vehicles with speed <0.1m/s<0.1m/s, normalized by lane capacity; N-dimensional.
@@ Appendix B Ablation Study
To investigate the contribution of each core component in the proposed HiLight framework, an ablation study is conducted by progressively removing key modules and evaluating the resulting performance during training. As shown in the Episode–Reward curves in Figure, the complete HiLight model achieves the best overall performance, exhibiting both the fastest convergence speed and the most stable behavior after convergence.
@@
We also examine the impact of pretraining the Meta-Policy prior to full model training, as shown in Figure. The Episode–Reward curve shows a smoother and faster start, without the initial reward drop observed in the baseline. Pretraining allows the Meta-Policy to generate more informative sub-goals in early training, which effectively guides the Sub-Policy toward better optimization directions and improves exploration efficiency. Although the learning slope becomes gentler in later stages, convergence speed is slightly improved, and the final performance remains on par with the non-pretrained HiLight.
@@ Appendix D Visualizations of the scenarios
Figure, Figure, and Figure illustrate the traffic scenarios used in our experiments.
@@ Appendix E Hyper-parameter Settings
Entropy Coefficient $0.01$ Number of Actions 88
@@
Discount Factor $0.99$ Yellow Light Duration (s) 55
@@
Learning Rate 3e43e-4 SUMO Start Port $15000$
@@ Appendix H Broad Impacts
Meta-Policy Transformer Encoder Extract global and regional spatiotemporal features across subregions Transformer Encoder: 3 layers, 2 heads, hidden size 4, feedforward dim = 165, no layer norm $(20,4,4)$: 20 time steps, 4 regions, 4-dim observation per region Adds a class token and positional encoding $\rightarrow$ Transformer encoder $\rightarrow$ output $(5,20,4)$ $\rightarrow$ extract $(1,1,4)$ as global feature; transform remaining $(4,20,4)$ $\rightarrow$ $(1,20,16)$
Meta-Policy MultiLayer LSTM Capture global temporal patterns across all regions 4-layer LSTM + FC layer $(1,20,16)$: sequence of flattened regional features LSTM (hidden size = 256) $\rightarrow$ use last-step output $\rightarrow$ FC layer $\rightarrow$ $(1,1,16)$ $\rightarrow$ reshape to $(1,4,4)$
@@
Feature Fusion Combine global and local features as input for shared policy network Broadcast + Concatenation Global: $(1,1,4)$, Local: $(1,16,330)$ Broadcast global feature to $(1,16,4)$ $\rightarrow$ concatenate with local features $\rightarrow$ fused tensor $(1,16,334)$
Shared Head for Critic & Actor Project fused traffic features into shared latent representation FC layers: $334\rightarrow 256\rightarrow 128\rightarrow 114$ (with dropout) $(1,16,334)$: concatenated per-node features Three-layer MLP with ReLU and dropout $\rightarrow$ output $(1,16,114)$ $\rightarrow$ used by actor and critic
@@
Actor Model Output traffic signal control actions 2-layer MLP with ReLU activation $(16,114)$: shared feature per signal $114\rightarrow 57\rightarrow 8\Rightarrow$ action logits $(16,)$